3. The following particulars relate to a symmetrical tangent cam operating a roller follower :-

Least radius = 36 mm, nose radius = 30 mm, roller radius = 21 mm, distance between cam

shaft and nose centre = 28 mm, angle of action of cam = 150°, cam shaft speed = 750 r.p.m.

Assuming that there is no dwell between ascent and descent, determine the lift of the valve

and the acceleration of the follower at a point where straight flank merges into the circular

nose.


1
Expert's answer
2021-07-08T04:02:03-0400

We have,

Total lift + r1 = PO + r2

36 + 30 = d – 5 d

= 54 mm

"R = (r1^2- r2^2 =d^2-2r1.d.cos \alpha)/ 2(r1-r2-dcos\alpha"

"R= (36^2-5^2 =36^2-2*30*36*cos150)\/ 2(30-5-45cos150)"

R= 72.42mm

therefore PQ = R-r2=82.41-5 = 87.41mm

QO = R-ri = 82.41-30 = 62.41mm